a toy rocket has a total mass of 0.1 kg including fuel  and  fuel’s mass is 0.02 kg which burns in 3 seconds. starting from rest on horizontal smooth track the rocket gets a velocity of 20 after the fuel is burned out. find the thrust on rocket

Grade:9

1 Answers

Rock Guitarian
8 years ago
thrust=mg 
=mass * gravity
=(mass of the rocket +mass of the fuel)* gravity
=0.12*10=1.2
1.2 – 0.12*20/3 =1.2 -0.8=0.4
ithink so this is correct once verify